OpenStudy (anonymous):

Find the value of the variable if P is between J & K. JP = 3y+1 ; PK = 12y - 4 ; JK = 75 Choose one answer. a. 5.2 b. 8.7 c. 3.9 d. 4.3

OpenStudy (anonymous):

First draw the diagram. or picture the problem. hint: P is between J and K whichn means. : J...........P...........K Given JK = 75 : <-------------75-------> Now, to find out the value of variable (y): distance JP = 3y+1 :call it eq1 distance PK = 12y-4 :call it eq2 distance J to P and from P to K = 75 that is: JK = JP+PK=75 :substitude the value of JP and PK from eq1 and e2 respectively, we get: = (3y+1) + (12y-4) = 3y+1+12y-4 = 15y-3 75 = 15y-3 :Take 3 common 25 = 5y-1 or 5y = 26 or y = 26/5 = 5.2 = (a.) <= answer